# Dissipativity-Based Synthesis of Distributed Control and Communication Topology Co-Design for AC Microgrids

### v3

@@ Abstract
This paper introduces a dissipativity-based framework for the joint design of distributed controllers and communication topologies in AC microgrids (MGs), providing robust performance guarantees for voltage regulation, frequency synchronization, and proportional power sharing across distributed generators (DGs). The closed-loop AC MG is represented as a networked system in which DGs, distribution lines, and loads function as interconnected subsystems linked through cyber-physical networks. Each DG utilizes a three-layer hierarchical control structure: a steady-state controller for operating point configuration, a local feedback controller for voltage tracking, and a distributed droop-free controller implementing normalized power consensus for frequency coordination and proportional power distribution. The operating point design is formulated as an optimization problem. Leveraging dissipativity theory, we derive necessary and sufficient subsystem dissipativity conditions. The global co-design is then cast as a convex linear matrix inequality (LMI) optimization that jointly determines distributed controller parameters and sparse communication architecture while managing the highly nonlinear, coupled dq-frame dynamics characteristic of AC systems.
@@ V-B3 Load Subsystems
To guarantee the existence of an equilibrium point satisfying the voltage regulation, frequency synchronization, and power sharing objectives, the reference voltages $V_{i,ref}$ and power sharing coefficients $P_{s}$ and $Q_{s}$ must be feasible solutions of the following optimization problem:
$\min_{V_{r},P_{s},Q_{s}}$ $\alpha_{V}\|V_{r}-\bar{V}_{r}\|^{2}+\alpha_{P}P_{s}+\alpha_{Q}Q_{s}$ (49)
s.t.: $V_{\min}\leq|V_{r}^{dq}|\leq V_{\max},0\leq P_{s}\leq 1,0\leq Q_{s}\leq 1,$
$P_{\max}\mathbf{1}_{N}P_{s}=\text{diag}(V_{r}^{d})I_{E}^{d}+\text{diag}(V_{r}^{q})I_{E}^{q},$
$Q_{\max}\mathbf{1}_{N}Q_{s}=\text{diag}(V_{r}^{q})I_{E}^{d}-\text{diag}(V_{r}^{d})I_{E}^{q},$
where $V_{r}=[V_{i,ref}^{dq}]_{i\in\mathbb{N}_{N}}$, $V_{r}^{d}$ and $V_{r}^{q}$ are the $d$ and $q$ components of $V_{r}$, $I_{E}^{dq}$ is determined from the network equilibrium equations (43d)-(43e) as a function of $V_{r}$, $P_{\max}\triangleq\text{diag}([P_{i}^{\max}]_{i\in\mathbb{N}_{N}})$, $Q_{\max}\triangleq\text{diag}([Q_{i}^{\max}]_{i\in\mathbb{N}_{N}})$, $\bar{V}_{r}$ denotes the desired reference voltage values, and $\alpha_{V}>0$, $\alpha_{P}>0$, and $\alpha_{Q}>0$ are normalizing cost coefficients.
Overall, this formulation ensures proper system operation through multiple aspects. The equality constraints guarantee that both active and reactive proportional power sharing objectives are achieved across all DGs simultaneously. The reference voltage magnitude bounds maintain system operation within safe and efficient limits. Furthermore, the constraints on $P_{s}$ and $Q_{s}$ ensure that the power sharing coefficients remain properly normalized for practical implementations.
The AC MG requires two sets of coupled equality constraints, one for active power and one for reactive power, due to the dual power flow objectives inherent to AC systems. These equality constraints couple the power sharing requirements with the network power flow equations (5), where the line injection currents $I_{E}^{dq}$ depend on the reference voltages through the network equilibrium relationships (43d)-(43e). This introduces additional complexity in determining a feasible set of reference voltages and power sharing coefficients that simultaneously satisfy voltage regulation, frequency synchronization, and proportional power sharing objectives.
@@ VI-B Local Controller Synthesis
VII Conclusion

### v4

@@ Abstract
This paper introduces a dissipativity-based framework for the joint design of distributed controllers and communication topologies in AC microgrids (MGs), providing robust performance guarantees for voltage regulation, frequency synchronization, and proportional power sharing across distributed generators (DGs). The closed-loop AC MG is represented as a networked system in which DGs, distribution lines, and loads function as interconnected subsystems linked through cyber-physical networks. Each DG utilizes a three-layer hierarchical control structure: a steady-state controller for operating point configuration, a local feedback controller for voltage tracking, and a distributed droop-free controller implementing normalized power consensus for frequency coordination and proportional power distribution. The operating point design is formulated as an optimization problem. Leveraging dissipativity theory, we derive necessary and sufficient subsystem dissipativity conditions. The global co-design is then cast as a convex linear matrix inequality (LMI) optimization that jointly determines distributed controller parameters and sparse communication architecture while managing the highly nonlinear, coupled dq-frame dynamics characteristic of AC systems. Simulation results from an islanded AC MG in a MATLAB/Simulink environment verify that the proposed framework achieves robust voltage regulation, frequency synchronization, and proportional power sharing through the optimized communication topology.
@@ V-B3 Load Subsystems
To guarantee the existence of an equilibrium point satisfying voltage regulation, frequency synchronization, and proportional power sharing objectives, the reference voltages $V_{r}\in\mathbb{R}^{2N}$ and active power sharing coefficients $\mathbf{P}_{s}=\begin{bmatrix}P_{s,1},\ldots,P_{s,N}\end{bmatrix}^{\top}\in\mathbb{R}^{N}$ must be feasible solutions of the following optimization problem:
$\min_{V_{r},\mathbf{P}_{s}}\alpha_{V}\|V_{r}-\bar{V}_{r}\|^{2}+\sum_{i=1}^{N}\left[\alpha_{P}(P_{s,i}-\bar{P}_{s})^{2}+\alpha_{Q}(Q_{s,i}-\bar{Q}_{s})^{2}\right]$ (49)
$\mbox{s.t.:}\ V_{\min}\leq|V_{i,ref}^{dq}|\leq V_{\max},\quad 0\leq P_{s,i}\leq 1,\ \forall i\in\mathcal{N}_{N},$
$P_{i}=P_{\max,i}\cdot P_{s,i},\ \forall i\in\mathcal{N}_{N},$
$I_{E}^{dq}=(M^{dq})^{-1}\begin{bmatrix}(Z^{dq})^{-1}\mathcal{B}_{D}^{\top}V_{r}-(Z^{dq})^{-1}\mathcal{B}_{L}^{\top}(\check{Y}^{dq})^{-1}\check{\bar{I}}_{L}^{dq}\end{bmatrix}$,
where $V_{r}=\begin{bmatrix}V_{1,ref}^{dq},...,V_{N,ref}^{dq}\end{bmatrix}^{\top}\in\mathbb{R}^{2N}$ stacks the DG reference voltages with $V_{i,ref}^{dq}=\begin{bmatrix}V_{i,ref}^{d}&V_{i,ref}^{q}\end{bmatrix}^{\top}$, $\bar{V}_{r}=\begin{bmatrix}\bar{V}_{d}&\mathbf{0}\end{bmatrix}^{\top}$ is the nominal reference where $\bar{V}_{d}=\begin{bmatrix}V_{d,nom,1},...,V_{d,nom,N}\end{bmatrix}^{\top}$ and $q$-axis references are zero. $P_{i}=V_{i,ref}^{d}I_{i}^{d}+V_{i,ref}^{q}I_{i}^{q}$ is the active power generated by $\Sigma_{i}^{DG}$ at equilibrium, $P_{s,i}=P_{i}/P_{\max,i}\in[0,1]$ is the normalized active power coefficient for $\Sigma_{i}^{DG}$, $Q_{s,i}=Q_{i}/Q_{\max,i}$ is the normalized reactive power coefficient where $Q_{i}=V_{i,ref}^{q}I_{i}^{d}-V_{i,ref}^{d}I_{i}^{q}$, $\bar{P}_{s}=\frac{1}{N}\sum_{i=1}^{N}P_{s,i}$ and $\bar{Q}_{s}=\frac{1}{N}\sum_{i=1}^{N}Q_{s,i}$ are the average normalized coefficients, $I_{i}^{dq}=\begin{bmatrix}\mathcal{B}_{D}I_{E}^{d}&\mathcal{B}_{D}I_{E}^{q}\end{bmatrix}_{i}^{\top}\in\mathbb{R}^{2}$ is the current injection at $\Sigma_{i}^{DG}$ obtained from the network equations, $I_{E}^{dq}\in\mathbb{R}^{2L}$ are the line equilibrium currents with $d$- and $q$- axis components $I_{E}^{d},I_{E}^{q}\in\mathbb{R}^{L}$, $M^{dq}\triangleq\mathbf{I}+(Z^{dq})^{-1}\mathcal{B}_{L}^{\top}(\check{Y}^{dq})^{-1}\mathcal{B}_{L}$ is the network coupling matrix, $\mathcal{B}_{D}\in\mathbb{R}^{N\times L}$, $\mathcal{B}_{L}\in\mathbb{R}^{M\times L}$ are the DG-to-line and load-to-line incidence matrices, $Z^{dq}\triangleq\text{diag}([R_{l}+\mathrm{i}\omega_{0}L_{l}]_{l\in\mathbb{N}_{L}})$ is the line impedance matrix, $\check{Y}^{dq}\triangleq\text{diag}([Y_{Lm}+\mathrm{i}\omega_{0}C_{tm}]_{m\in\mathbb{N}_{M}})$ is the load admittance matrix, $\check{\bar{I}}_{L}^{dq}\in\mathbb{R}^{2M}$ are the constant load currents, and $\alpha_{V},\alpha_{P},\alpha_{Q}>0$ are weighting coefficients.
@@ VI-B Local Controller Synthesis
VII Simulation Results
To demonstrate the effectiveness of the dissipativity-based distributed control and communication topology co-design framework, we simulated an islanded AC MG in MATLAB. The AC MG comprises $N=4$ DGs and $L=8$ lines, with each DG supporting a local load. The nominal voltage amplitude was set to $V_{\rm ref}=311.13$ V (corresponding to $220$ V RMS), and the nominal grid frequency was set to $\omega_{0}=2\pi\times 60$ rad/s. Each DG is modeled with an LC filter and a ZIP load, with parameters generated using a random seed. The physical topology was generated using a minimum spanning tree (MST)-based algorithm with a connectivity threshold of $0.4$, resulting in $8$ physical lines connecting the $4$ DGs and $2$ global loads, as shown in Fig. 7(a). In the co-design, the $\mathcal{L}_{2}$-gain bound was selected as $\bar{\gamma}=1000$.
The local controllers achieve IF-OFP passivity indices of $\nu_{i}=-0.97$ and $\rho_{i}=1$ for all DGs, confirming that the dissipativity conditions of Th. 4 are satisfied. We evaluated two co-design variants to examine the trade-off between communication and closed-loop performance: (i) a hard graph constraint that restricts the communication topology $\mathcal{G}^{c}$ to exactly match the physical DG-to-DG adjacency, resulting in $4$ bidirectional communication links, as shown in Fig. 7(b); and (ii) a soft graph constraint that penalizes communication links via a $\ell_{1}$ sparsity term in the LMI cost function (see Th. 3), allowing the co-design to determine a sparser topology, as shown in Fig. 7(c). The resulting soft-constrained topology contains only $2$ undirected links, representing a $50\%$ reduction in communication links relative to the hard graph case and a communication density of $33\%$ compared to the $67\%$ density of the physical topology.
Figure 8 summarizes the closed-loop performance of the proposed framework. As shown in Fig. 8(a), all DG voltages converge to the reference value $V_{\rm ref}=311$ V from a cold start (i.e., starting from zero initial conditions), confirming effective voltage regulation with zero steady-state error across all DGs. Figure 8(b) shows that all DG frequencies track the nominal reference value of $60$ Hz, demonstrating successful frequency synchronization. As shown in Fig. 8(c), all DGs converge to the same normalized active power level of $P_{i}/P_{i}^{\rm max}=0.6$, achieving proportional active power sharing. Figure 8(d) shows that the normalized reactive power levels also converge, confirming effective reactive power sharing. Notably, the performance under the soft graph constraint is identical to that under the hard graph constraint across all four metrics, demonstrating that the proposed co-design framework can reduce communication infrastructure cost without sacrificing closed-loop performance.
Figure 7: The AC MG topology with $N=4$ DGs, $L=8$ lines, and $M=2$ global loads: (a) physical, (b) hard graph communication matching physical adjacency, and (c) soft graph communication with reduced overhead.
Figure 8: The closed-loop performance of the proposed dissipativity-based distributed controller in the AC MG with 4 DGs and 8 lines: (a) DG voltage magnitudes, (b) DG frequencies, (c) normalized active power sharing, and (d) normalized reactive power sharing.
VIII Conclusion